OpenStudy (anonymous):

I would appreciate an actual answer without swearing please. Which statement about the simplified binomial expansion of (a – b)n, where n is a positive integer, is true? All terms of the expansion are positive. All terms in the expansion are negative. The first term in the expansion is positive. The first term in the expansion is negative.

OpenStudy (anonymous):

@Heather-Bates Excuse Me?

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

Swearing is banned in OS. When you swear instead of the actual word it changes it to something else random.

OpenStudy (texaschic101):

I think it is : the first term in the expansion is positive

OpenStudy (anonymous):

Hey, Texas, why d'you think that's the case? I don't really understand how I could come to any particular conclusion on this question.

OpenStudy (texaschic101):

sub in a positive integer for n : (a - b)^2 = a^2 - 2ab + b^2 (a - b)^4 = a^4 -4a^3b + 6a^2b^2 - 4ab^3 + b^4 (a-b)^6 = it starts with a positive a^6 Or am I misunderstanding the question...is the n supposed to be ^n

OpenStudy (texaschic101):

or is it supposed to be (a - b)2 meaning 2a - 2b ??

OpenStudy (anonymous):

Hmmm...That's curious.

OpenStudy (texaschic101):

is the n supposed to be an exponent ?

OpenStudy (anonymous):

If n was an exponent, we'd have to deal with (a-b)(a-b), which gives us 3 terms instead of two. We know that this is a binomial, which means we can only have 2 terms, and so n must be a normal number and not an exponent.

OpenStudy (texaschic101):

ok... (a - b)2 = 2a - 2b (a - b)4 = 4a - 4b (a - b)6 = 6a - 6b first term is positive

OpenStudy (anonymous):

Yeah, you're right. Texaschic is right.

OpenStudy (texaschic101):

:)...confirmation is always good

OpenStudy (anonymous):

I just don't like it when more and more people come on and offer different solutions when the real answer sticks out like a sore thumb.

OpenStudy (anonymous):

@texaschic101 Thank you so much. :)

OpenStudy (texaschic101):

no problem....thats what I am here for